# On computability of equilibrium states

### v2

@@ 6. Non-uniqueness of equilibrium states
In this section, we establish Theorem E and then use it to prove Theorem D. We first recall a classical result in Computable Analysis.
Lemma 6.1 ([GHR10, Proposition 7]).
@@
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ be a continuous map. Then $\operatorname{supp}(\mu)\subseteq\Omega(T)$ for each $\mu\in\mathcal{M}_{T}(X)$.
$\Omega(T)$ is the non-wandering set of $T$ defined in (2.1).
@@
Consider $\mu\in\mathcal{M}_{T}(X)$. By [Wa82, Theorem 5.6 (i)], $\Omega(T)$ is closed. Moreover, it follows from [Wa82, Theorem 6.15 (i)] that $\mu(\Omega(T))=1$. Therefore, $\operatorname{supp}(\mu)\subseteq\Omega(T)$. ∎
@@
$\mathcal{H}(\mu)\coloneqq\begin{cases}h_{\mu}(T)&\text{if }\mu\in\mathcal{M}_{T}(X);\\ 0&\text{if }\mu\notin\mathcal{M}_{T}(X)\end{cases}$
is upper-computable on $\mathcal{M}_{T}(X)$, and that $\varphi\colon X\rightarrow\mathbb{R}$ is a computable potential satisfying that the topological pressure $P(T,\varphi)$ is lower-computable. Then the set of equilibrium states of $T$ for $\varphi$ is recursively compact.
@@
By Corollary 3.20, the integral function $\Phi\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ given by $\Phi(\mu)\coloneqq\int\!\varphi\,\mathrm{d}\mu$, $\mu\in\mathcal{P}(X)$, is computable. Since $\mathcal{H}(\mu)$ is upper-computable on $\mathcal{M}_{T}(X)$, then the measure-theoretic pressure function $\mathcal{P}(\mu)\coloneqq\mathcal{H}(\mu)+\Phi(\mu)$ is upper-computable on $\mathcal{M}_{T}(X)$. Since $P(T,\varphi)$ is lower-computable, then there exists a uniformly computable sequence of rational points $\{r_{n}\}_{n\in\mathbb{N}}$ which is increasing and converges to $P(T,\varphi)$. By Definition 3.12, there exists a uniformly sequence $\{U_{n}\}_{n\in\mathbb{N}}$ of lower-computable open sets satisfying that $\mathcal{P}^{-1}(-\infty,r_{n})\cap\mathcal{M}_{T}(X)=U_{n}\cap\mathcal{M}_{T}(X)$ for each $n\in\mathbb{N}$. Hence,
(6.1) $\mathcal{M}_{T}(X)\setminus\mathcal{P}^{-1}(-\infty,P(T,\varphi))$ $=\mathcal{M}_{T}(X)\setminus\Bigl{(}\bigcup_{n\in\mathbb{N}}\mathcal{P}^{-1}(-\infty,r_{n})\cap\mathcal{M}_{T}(X)\Bigr{)}$
$=\Bigl{(}\bigcup_{n\in\mathbb{N}}U_{n}\Bigr{)}^{c}\cap\mathcal{M}_{T}(X).$
By Definition 2.4, this set consists of all equilibrium states of $T$ for $\varphi$. Moreover, by Propositions 3.7 and 3.15, it follows from (6.1) that this set is recursively compact. ∎
Now we are ready to prove Theorem E.
@@
We prove Theorem E by contradiction and suppose that $\mu$ is the unique equilibrium state of $T$ for $\varphi$. Then by Proposition 3.15 (i), it follows from Lemma 6.3 that $\mu$ is computable, which would contradict with the existence of non-computable equilibrium state. Noting that every upper-computable function is upper semi-continuous, one can see that $h_{\mu}(T)$ is upper semi-continuous on $\mu\in\mathcal{M}_{T}(X)$. Hence, by [PU10, Proposition 3.5.3], there exists at least one equilibrium state of $T$ for $\varphi$. Thus, statement (ii) holds. Therefore, there are at least two equilibrium states of $T$ for $\varphi$. ∎
As an immediate consequence, we establish Theorem D.
Proof of Theorem D.
Assume that $T\colon X\rightarrow X$ is a computable map with zero topological entropy satisfying that there exists no computable point in $\Omega(T)$. Then by [PU10, Proposition 3.5.3], there exists at least one equilibrium state of $T$ for $\varphi$. Suppose that $\mu$ is an equilibrium state of $T$ for $\varphi$. Then by Lemma 6.2, there exists no computable points in $\operatorname{supp}(\mu)$. It follows from Lemma 6.1 that $\mu$ is non-computable. Therefore, we can derive Theorem E from Theorem D. ∎
@@ 7.1. Proof of Theorem F
In this subsection, we demonstrate that the computable homeomorphism $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ constructed in [GHR10, Subsection 4.1] satisfies the statements in Theorem F for each potential $\varphi\in C(\mathbb{S}^{1})$.
@@
Recall that $T$ constructed in [GHR10, Subsection 4.1] is a computable homeomorphism of the unit circle $\mathbb{S}^{1}$ whose non-wandering set $\Omega(T)$ contains no computable points. Since $T$ is a homeomorphism of the circle, it follows from [Wa82, Theorem 7.14] that the topological entropy of $T$ is zero, which implies that $T$ satisfies statement (i) of Theorem F. Suppose that $\mu$ is an equilibrium state of $T$ for $\varphi$. Then by Lemma 6.2, there exists no computable points in $\operatorname{supp}(\mu)$. It follows from Lemma 6.1 that $\mu$ is non-computable. Therefore $T$ satisfies statement (ii) of Theorem F. ∎
@@ 7.2. Asymptotic h-expansive
In this subsection, we introduce additional relevant notions and results from ergodic theory. We use them to construct computable systems with arbitrarily high topological entropy whose equilibrium states are all non-computable. We start by recalling the following definition from [Dow11, Remark 6.1.7].
@@
By the Lebesgue Covering Lemma, it is clear that for a compact metric space, refining sequences of open covers always exist. The topological tail entropy was first introduced by M. Misiurewicz under the name “topological conditional entropy” [Mis73, Mis76]. We use the terminology in [Dow11] (see [Dow11, Remark 6.3.18] for reference).
@@
Let $(X,d)$ be a compact metric space and $g\colon X\rightarrow X$ a continuous map. The topological conditional entropy $h(g|\lambda)$ of $g$ given $\lambda$, for some open cover $\lambda$, is defined as
@@
where $\{\xi_{l}\}_{l\in\mathbb{N}_{0}}$ is an arbitrary refining sequence of open covers, and for each pair of open covers $\xi$ and $\eta$,
@@
is the logarithm of the minimal number of sets from $\xi$ sufficient to cover any set in $\eta$. The topological tail entropy $h^{*}(g)$ of $g$ is defined by
@@
where $\{\xi_{l}\}_{l\in\mathbb{N}_{0}}$ and $\{\eta_{m}\}_{m\in\mathbb{N}_{0}}$ are two arbitrary refining sequences of open covers, and $H$ is as defined in (7.2).
The concept of $h$-expansiveness was introduced by R. Bowen in [Bo72]. We use the formulation in [Mis76] (see also [Dow11]).
@@
A continuous map $g\colon X\rightarrow X$ on a compact metric space $(X,d)$ is called $h$-expansive if there exists a finite open cover $\lambda$ of $X$ such that $h(g|\lambda)=0$.
Remarks 7.4.
As an equivalent definition of $h$-expansiveness, the map $g\colon X\rightarrow X$ is called $h$-expansive if there exists $\epsilon>0$ such that
$h_{\operatorname{top}}\bigl{(}g|_{{\Phi}_{\epsilon,g}(x)}\bigr{)}=0$
for each $x\in X$, where
@@
A weaker property was introduced by M. Misiurewicz in [Mis73] (see also [Mis76, Dow11]).
@@
Remarks 7.6.
The topological entropy of $g$ is $h_{\operatorname{top}}(g)=h(g|\{X\})$, where $\{X\}$ is the open cover of $X$ consisting of only one open set $X$. It also follows from Definition 7.1, for open covers $\iota$ and $\eta$ of $X$, one can see $h(g|\iota)\leqslant h(g|\eta)$ if $\iota$ is a refinement of $\eta$. Hence, we have $h^{*}(g)\leqslant h_{\operatorname{top}}(g)$.
@@
Let $(X,f)$ be a cascade. If $f$ is asymptotically $h$-expansive, then for each $\varphi\in C(X)$, there exists at least one equilibrium state of $f$.
@@
Assume $d\in\mathbb{N}$ and let $T_{d}\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ be given by $T_{d}(x)\coloneqq dx\pmod{1}$ for $x\in\mathbb{S}$. Let $\widehat{T}\coloneqq T\times T_{d}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$ be defined as in (7.4) and $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$ be a continuous potential. Then statements (i) and (ii) in Theorem G hold for $\widehat{T}$ and $\widehat{\varphi}$, and $h_{\operatorname{top}}\bigl{(}\widehat{T}\bigr{)}=\log d$.
@@
First, we demonstrate that $\widehat{T}$ satisfies statement (i) in Theorem G. Note that for the map $T_{d}$ and sufficiently small $\epsilon>0$, the set $\Phi_{\epsilon,T_{d}}(x)=\{x\}$ for each $x\in\mathbb{S}^{1}$ (see (7.3)). Then by Remark 7.4, $T_{d}$ is $h$-expansive, and thus $h^{*}(T_{d})=0$. Noting that $h_{\operatorname{top}}(T)=0$, it follows from Remark 7.6 that $h^{*}(T)\leqslant h_{\operatorname{top}}(T)=0$, hence, is equal to 00. Hence by Proposition 7.7, $h^{*}\bigl{(}\widehat{T}\bigr{)}=h^{*}(T)+h^{*}(T_{d})=0,$ which implies that $\widehat{T}$ is asymptotically $h$-expansive. By Proposition 7.8, for each continuous potential $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$, there exists at least one equilibrium state of $\widehat{T}$. Finally, we argue that there exists no computable $\widehat{T}$-invariant Borel probability measure by contradiction and assume that $\widehat{\mu}\in\mathcal{M}_{\widehat{T}}(\mathbb{S}^{1}\times\mathbb{S}^{1})$ is computable. Then $\mu$ given by $\mu(A)\coloneqq\widehat{\mu}(A\times\mathbb{S}^{1})$ for each Borel set $A\subseteq\mathbb{S}^{1}$ is a computable $T$-invariant probability measure. This is impossible by Theorem F (ii). Thus $\widehat{T}$ satisfies statement (ii) in Theorem G. ∎

### v3

@@ 6. Non-uniqueness of equilibrium states
In this section, we establish Theorem D and then use it to prove Theorem E.
Lemma 6.1 ([GHR11, Proposition 7]).
@@
Let $(X,\rho)$ be a compact metric space, and $T\colon X\rightarrow X$ be a continuous map. Then for each $\mu\in\mathcal{M}(X,T)$, $\operatorname{supp}(\mu)$ is contained in the non-wandering set $\Omega(T)$ of $T$.
@@
Consider $\mu\in\mathcal{M}(X,T)$. By [Wa82, Theorem 6.15 (i)], we have $\mu(\Omega(T))=1$. Note that by [Wa82, Theorem 5.6 (i)], $\Omega(T)$ is closed. Then we obtain that $\operatorname{supp}(\mu)\subseteq\Omega(T)$. ∎
@@
$\mathcal{H}(\mu)\coloneqq\begin{cases}h_{\mu}(T)&\text{if }\mu\in\mathcal{M}(X,T);\\ 0&\text{if }\mu\notin\mathcal{M}(X,T)\end{cases}$
is upper-computable on $\mathcal{M}(X,T)$, and that $\varphi\colon X\rightarrow\mathbb{R}$ is a computable function satisfying that the topological pressure $P(T,\varphi)$ is lower-computable. Then the set $\mathcal{E}(T,\varphi)$ of equilibrium states for $T$ and $\varphi$ is recursively compact.
@@
By Corollary 3.20, the integral function $\Phi\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ given by $\Phi(\mu)\coloneqq\int\!\varphi\,\mathrm{d}\mu$, $\mu\in\mathcal{P}(X)$, is computable. Since $\mathcal{H}(\mu)$ is upper-computable on $\mathcal{M}(X,T)$, the measure-theoretic pressure function $\mathcal{P}(\mu)\coloneqq\mathcal{H}(\mu)+\Phi(\mu)$ is upper-computable on $\mathcal{M}(X,T)$.
Since $P(T,\varphi)$ is lower-computable, there exists a uniformly computable sequence of rational points $\{r_{n}\}_{n\in\mathbb{N}}$ which is increasing and converges to $P(T,\varphi)$. By Definition 3.12, there exists a uniformly computable sequence $\{U_{n}\}_{n\in\mathbb{N}}$ of lower-computable open sets with $\mathcal{P}^{-1}(-\infty,r_{n})\cap\mathcal{M}(X,T)=U_{n}\cap\mathcal{M}(X,T)$ for each $n\in\mathbb{N}$. Hence, by Definition 2.5, we have
(6.1) $\mathcal{E}(T,\varphi)$ $=\mathcal{M}(X,T)\smallsetminus\mathcal{P}^{-1}(-\infty,P(T,\varphi))$
$=\mathcal{M}(X,T)\smallsetminus\Bigl{(}\bigcup_{n\in\mathbb{N}}\bigl{(}\mathcal{P}^{-1}(-\infty,r_{n})\cap\mathcal{M}(X,T)\bigr{)}\Bigr{)}=\mathcal{M}(X,T)\cap\Bigl{(}\bigcup_{n\in\mathbb{N}}U_{n}\Bigr{)}^{c}.$
Moreover, by Propositions 3.7 and 3.15 (iii), it follows from (6.1) that the set $\mathcal{E}(T,\varphi)$ of all equilibrium states for $T$ and $\varphi$ is recursively compact. ∎
Now, we are ready to prove Theorem D.
Proof of Theorem D.
First, we show that there exists at least one equilibrium state for $T$ and $\varphi$. Indeed, by Definition 3.12, every upper-computable function is upper semi-continuous. Since $\mathcal{H}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ is upper-computable on $\mathcal{M}(X,T)$, $\mathcal{H}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ is upper semi-continuous on $\mu\in\mathcal{M}(X,T)$. Hence, by [PU10, Proposition 3.5.3], there exists at least one equilibrium state for $T$ and $\varphi$.
Now we prove Theorem D by contradiction and suppose that $\mu$ is the unique equilibrium state for $T$ and $\varphi$. By Lemma 6.3, the singleton $\{\mu\}$ is a recursively compact set. Then by Proposition 3.15 (i), $\mu$ is computable, which would contradict the existence of a non-computable equilibrium state. Therefore, there are at least two equilibrium states for $T$ and $\varphi$. This completes the proof of Theorem D. ∎
As an immediate consequence, we establish Theorem E.
@@
Assume that $T\colon X\rightarrow X$ is a computable map with zero topological entropy satisfying that there exists no computable point in $\Omega(T)$. Then by [PU10, Proposition 3.5.3], there exists at least one equilibrium state for $T$ and $\varphi$, say $\mu$. Then by Lemma 6.2, there exists no computable point in $\operatorname{supp}(\mu)$. It follows from Lemma 6.1 that $\mu$ is non-computable. Therefore, by Theorem D, there exists at least two equilibrium states for $T$ and $\varphi$. ∎
@@ 7.1. Proof of Theorem F
In this subsection, we demonstrate that the computable homeomorphism $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ constructed in [GHR11, Subsection 4.1] satisfies the statements (i) and (ii) in Theorem F for each function $\varphi\in C\bigl{(}\mathbb{S}^{1}\bigr{)}$.
@@
Recall that the homeomorphism $T$, which is constructed in [GHR11, Subsection 4.1], is a computable homeomorphism of the unit circle $\mathbb{S}^{1}$ whose non-wandering set $\Omega(T)$ contains no computable points.
Since $T$ is a homeomorphism of the circle, by [Wa82, Theorem 7.14], we have $h_{\operatorname{top}}(T)=0$. Then, by [PU10, Proposition 3.5.3], there exists at least one equilibrium state for $T$ and $\varphi$. Hence, $T$ satisfies statement (i) in Theorem F.
Suppose that $\mu$ is an equilibrium state for $T$ and $\varphi$. Then by Lemma 6.2, there exists no computable point in $\operatorname{supp}(\mu)$. It follows from Lemma 6.1 that $\mu$ is non-computable. Therefore, $T$ satisfies statement (ii) in Theorem F. ∎
@@ 7.2. Asymptotic h-expansive
In this subsection, we introduce some additional relevant notions and results from ergodic theory. We use them to construct computable systems with arbitrarily high topological entropy whose equilibrium states are all non-computable. We begin by recalling the following definition from [Dow11, Remark 6.1.7].
@@
By Lebesgue’s covering lemma, it is clear that for each compact metric space, refining sequences of open covers always exist.
The topological tail entropy was first introduced by Misiurewicz under the name “topological conditional entropy” [Mis73, Mis76]. In our paper, we use the terminology in [Dow11] (see [Dow11, Remark 6.3.18] for reference).
@@
Let $(X,d)$ be a compact metric space, and $g\colon X\rightarrow X$ be a continuous map. The topological conditional entropy $h(g|\lambda)$ of $g$ for some open cover $\lambda$, is defined by
@@
where $\{\xi_{l}\}_{l\in\mathbb{N}_{0}}$ is an arbitrary refining sequence of open covers. For each pair of open covers $\xi$ and $\eta$,
@@
is the logarithm of the minimal number of sets from $\xi$ sufficient to cover any set in $\eta$.
The topological tail entropy $h^{*}(g)$ of $g$ is defined by
@@
where $\{\xi_{l}\}_{l\in\mathbb{N}_{0}}$ and $\{\eta_{m}\}_{m\in\mathbb{N}_{0}}$ are two arbitrary refining sequences of open covers, and $H$ is defined in (7.2).
The concept of $h$-expansiveness was introduced by Bowen in [Bo72]. We use the formulation in [Mis76] (see also [Dow11]).
@@
A continuous map $g\colon X\rightarrow X$ on a compact metric space $(X,d)$ is $h$-expansive if there exists a finite open cover $\lambda$ of $X$ such that $h(g|\lambda)=0$.
Remark 7.4.
Bowen gives an equivalent definition of $h$-expansiveness in [Bo72] as follows: the map $g\colon X\rightarrow X$ is called $h$-expansive if there exists $\epsilon>0$ with $h_{\operatorname{top}}\bigl{(}g|_{{\Phi}_{\epsilon,g}(x)}\bigr{)}=0$ for each $x\in X$, where
@@
A weaker property was introduced by Misiurewicz in [Mis73] (see also [Mis76, Dow11]).
@@
Remark 7.6.
The topological entropy of $g$ is $h_{\operatorname{top}}(g)=h(g|\{X\})$, where $\{X\}$ is the open cover of $X$ consisting of only one open set $X$. It also follows from Definition 7.1 that, for each pair of open covers $\iota$ and $\eta$ of $X$, we have that $h(g|\iota)\leqslant h(g|\eta)$ if $\iota$ is a refinement of $\eta$. Hence, we have $h^{*}(g)\leqslant h_{\operatorname{top}}(g)$.
@@
Let $(X,f)$ be a cascade. If $f$ is asymptotically $h$-expansive, then for each $\varphi\in C(X)$, there exists at least one equilibrium state for $f$ and $\varphi$.
@@
Assume that $d$ is an integer, and $T_{d}\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ is given by $T_{d}(x)\coloneqq dx\pmod{1}$ for $x\in\mathbb{S}^{1}$. Define $\widehat{T}\coloneqq T\times T_{d}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}\times\mathbb{S}^{1}$. Then statements (i) and (ii) in Theorem G hold for the map $\widehat{T}$ and each $\widehat{\varphi}\in C\bigl{(}\mathbb{S}^{1}\times\mathbb{S}^{1}\bigr{)}$, and $h_{\operatorname{top}}\bigl{(}\widehat{T}\bigr{)}=\log d$.
@@
First, we demonstrate that $\widehat{T}$ satisfies statement (i) in Theorem G. It is not hard to see that $T_{d}$ is a distance-expanding map. Then for sufficiently small $\epsilon>0$, the set $\Phi_{\epsilon,T_{d}}(x)=\{x\}$ for each $x\in\mathbb{S}^{1}$ (see (7.3)). Then by Remark 7.4, $T_{d}$ is $h$-expansive, and thus $h^{*}(T_{d})=0$. Noting that $h_{\operatorname{top}}(T)=0$, it follows from Remark 7.6 that $h^{*}(T)\leqslant h_{\operatorname{top}}(T)=0$, hence, is equal to 0. Hence by Proposition 7.7, $h^{*}\bigl{(}\widehat{T}\bigr{)}=h^{*}(T)+h^{*}(T_{d})=0,$ which implies that $\widehat{T}$ is asymptotically $h$-expansive. By Proposition 7.8, for each continuous potential $\widehat{\varphi}\colon\mathbb{S}^{1}\times\mathbb{S}^{1}\rightarrow\mathbb{R}$, there exists at least one equilibrium state for $\widehat{T}$ and $\widehat{\varphi}$.
Finally, we argue that there exists no computable $\widehat{T}-$invariant Borel probability measure by contradiction and assume that $\widehat{\mu}\in\mathcal{M}\bigl{(}\mathbb{S}^{1}\times\mathbb{S}^{1},\widehat{T}\bigr{)}$ is computable. Then $\mu$ given by $\mu(A)\coloneqq\widehat{\mu}\bigl{(}A\times\mathbb{S}^{1}\bigr{)}$ for each Borel set $A\subseteq\mathbb{S}^{1}$ is a computable $T$-invariant probability measure. This is impossible by Theorem F (ii). Thus $\widehat{T}$ satisfies statement (ii) in Theorem G. ∎